OpenStudy (anonymous):

you are studyingfor your final exam of the semester. Up to this point, you received 3 exam scores of 98%, 100%, and 64%. To receive a grade of A in the class, you must have an average exam score between 90% and 100% for all 4 exams including the final. Find the widest range of scores that you can get on the final exam in order to receive a grade of A for the class. Use interval notation. A.[91,100] B.[98,100] C.[88,97] Show work please.

OpenStudy (paxpolaris):

\[\Large 90 \le {98+100+64+x \over 4} \le 100\]

OpenStudy (paxpolaris):

\[\implies \large 90 \le {262+x \over 4}\]\[\implies \large 360 \le {262+x}\]\[\implies \large 360-262 \le x\] \[\implies \Large 98 \le x\]\[\implies \Large 98 \le x \le 100\]\[\implies \Large [98, 100]\]